OpenStudy (anonymous):

wat is - 2 3/10 x -4

OpenStudy (studygurl14):

-2 3/10 = -2.3 -2.3 • -4 = 9.2

OpenStudy (anonymous):

9.2

OpenStudy (studygurl14):

In fraction form, that's 9 1/5, or 19/2

OpenStudy (anonymous):

9.2